This is the first year Ralph is playing softball,he has been practicing per batting On his last swing the bat made an arc with a radius of 45 inches and swept through 250° of rotation Assuming the arc is circular what is the distance the tip of the bat travels to the nearest inch? How many feet is this rounded to the nearest foot?


1
Expert's answer
2021-10-27T12:56:34-0400



"L=\pi r \dfrac{\alpha}{180^{\circ}}\\\"

"r=45 \ inches" , "\alpha=250^{\circ}"


Then,

"L=\pi \times 45\times\dfrac{250^{\circ}}{180^{\circ}}=62.5\pi\ \ inches"


Now, In feet

"62.5\pi \ inches =196.35 \ inches"


We know that , 1 feet = 12 inches

So, 1 inch = 1/12 feet


hence, 196.35 inches = "196.35\times \dfrac{1}{12}=16.362\ \ feets\ \approx 16 foot"